¿ªÔÆÌåÓý

ctrl + shift + ? for shortcuts
© 2025 Groups.io

Tek 468 problem


 

Hi,
I have a Tek 468 analogue with digital display and storage. The problem: the scope works fine in analogue mode, however when digital mode is selected, the trace appears to be modulated at a higher frequency causing a blurred display. If the timebase x10 is used to expand the trace, it appears to be composed of a semi sinusoidal waveform.(Its not possible to determine the frequency as in the digital mode the timebase control doesnt seem to have much effect)and its not possible to re-create this effect in analogue mode. I have checked grounding connections, re-seated the i/cs etc all to no effect. My next step is to check the digital psu for excess ripple.
Has anyone come across this problem before or can anyone point me in the right direction?
Note: The digital functions seem to work ok,save, store etc.
Very often the 'scope will not 'POST (power on self test,) but will come on in analogue mode.
The digital LED display does work, and so do the cursors and they do measure, but the cursors are modulated in the same way as the trace.

Bill


Jim Popwell Jr
 

Hi Bill,
ripple? ?your on the right track I believe.

try checking the ripple output on all the power supplies. ?the caps in these older scopes go bad with age and the first sign is ripple on the DC supplies.
recapping the main electrolytics involved in the power supplies usually cures.
?jim


On Jul 10, 2012, at 2:38 AM, maninashed1 wrote:

?

Hi,
I have a Tek 468 analogue with digital display and storage. The problem: the scope works fine in analogue mode, however when digital mode is selected, the trace appears to be modulated at a higher frequency causing a blurred display. If the timebase x10 is used to expand the trace, it appears to be composed of a semi sinusoidal waveform.(Its not possible to determine the frequency as in the digital mode the timebase control doesnt seem to have much effect)and its not possible to re-create this effect in analogue mode. I have checked grounding connections, re-seated the i/cs etc all to no effect. My next step is to check the digital psu for excess ripple.
Has anyone come across this problem before or can anyone point me in the right direction?
Note: The digital functions seem to work ok,save, store etc.
Very often the 'scope will not 'POST (power on self test,) but will come on in analogue mode.
The digital LED display does work, and so do the cursors and they do measure, but the cursors are modulated in the same way as the trace.

Bill



 

Ive added some photos of the scope displaying the fault. These are in my photo folder under bill.valvetechnology.
Ive checked the waveforms and ripple from both power supplies against the manual and both the analogue and digital power supplies look fine.
Can anyone identify this fault? It has been suggested that it could be the clock signal being imposed on the waveform and the cursors in digital mode.
Thanks
Bill


 

Bill,

I'm wondering if that's noise from the D-to-A converter circuit, maybe because a low-pass filter cap is open. I'd suggest to check on the Storage Display Logic Board (A16), which is the top-most board. Check around op-amps U571A, U352 and U558. See if you've got that noise on the signals on pin 6 of those last 2 op-amps, and pin 1 of U571A. Oh, In between, there's a 250 kHz switching pulse on U523D. Hhhhuuummmm????

Jimmy

--- In TekScopes@..., "maninashed1" <bill.valvetechnology@...> wrote:



Ive added some photos of the scope displaying the fault. These are in my photo folder under bill.valvetechnology.
Ive checked the waveforms and ripple from both power supplies against the manual and both the analogue and digital power supplies look fine.
Can anyone identify this fault? It has been suggested that it could be the clock signal being imposed on the waveform and the cursors in digital mode.
Thanks
Bill


 

Hi Jimmy, and thank you for that information. Checked U571 - no problem, checked U558 op-amp faulty waveform present. In between is U352, a LF351 op-amp, this proved to be faulty and the source of the oscillation. New LF351 and scope now working fine again. Just needs a good service and thorough clean. Im going to keep this one, its a lovely bit of kit and is so useable for the things Im interested in. Thanks to everyone who took time to read this thread.
Bill

--- In TekScopes@..., "jtjewell83" <jtjewell83@...> wrote:

Bill,

I'm wondering if that's noise from the D-to-A converter circuit, maybe because a low-pass filter cap is open. I'd suggest to check on the Storage Display Logic Board (A16), which is the top-most board. Check around op-amps U571A, U352 and U558. See if you've got that noise on the signals on pin 6 of those last 2 op-amps, and pin 1 of U571A. Oh, In between, there's a 250 kHz switching pulse on U523D. Hhhhuuummmm????

Jimmy

--- In TekScopes@..., "maninashed1" <bill.valvetechnology@> wrote:



Ive added some photos of the scope displaying the fault. These are in my photo folder under bill.valvetechnology.
Ive checked the waveforms and ripple from both power supplies against the manual and both the analogue and digital power supplies look fine.
Can anyone identify this fault? It has been suggested that it could be the clock signal being imposed on the waveform and the cursors in digital mode.
Thanks
Bill


 

Great Bill!! Glad I could help.

I have 2 working 468s, and 2 parts units. The first working unit had bad PROMs when I got it, and several bad big electrolytic caps in the main power supply. The other working unit had a bad TTL clock oscillator in the microprocessor section, a bad U571 (if I remember correctly), and lots of minor problems like bad pots, broken knobs, etc. They both work and look great now.

When I get time, I enjoy fixing up old classic Tek and HP scopes. I use them from time-to-time. Last time I used the storage function on one of the 468s was troubleshooting on a PC's motherboard that had blown MOSFETs in the CPU power supply per bad electrolytics one the core supply's output.

Jimmy

--- In TekScopes@..., "maninashed1" <bill.valvetechnology@...> wrote:



Hi Jimmy, and thank you for that information. Checked U571 - no problem, checked U558 op-amp faulty waveform present. In between is U352, a LF351 op-amp, this proved to be faulty and the source of the oscillation. New LF351 and scope now working fine again. Just needs a good service and thorough clean. Im going to keep this one, its a lovely bit of kit and is so useable for the things Im interested in. Thanks to everyone who took time to read this thread.
Bill



--- In TekScopes@..., "jtjewell83" <jtjewell83@> wrote:

Bill,

I'm wondering if that's noise from the D-to-A converter circuit, maybe because a low-pass filter cap is open. I'd suggest to check on the Storage Display Logic Board (A16), which is the top-most board. Check around op-amps U571A, U352 and U558. See if you've got that noise on the signals on pin 6 of those last 2 op-amps, and pin 1 of U571A. Oh, In between, there's a 250 kHz switching pulse on U523D. Hhhhuuummmm????

Jimmy

--- In TekScopes@..., "maninashed1" <bill.valvetechnology@> wrote:



Ive added some photos of the scope displaying the fault. These are in my photo folder under bill.valvetechnology.
Ive checked the waveforms and ripple from both power supplies against the manual and both the analogue and digital power supplies look fine.
Can anyone identify this fault? It has been suggested that it could be the clock signal being imposed on the waveform and the cursors in digital mode.
Thanks
Bill